# biblio.bib

### v1

@@ 5.1 Test on a Triangle
plus the reversed arcs (Figure 1).
$x$$y$$z_{1}$$z_{3}$$z_{2}$$\gamma_{1}$$\gamma_{2}$$\gamma_{3}$
@@ 5.1.1 Hamiltonian Dependent on $s$
We begin evaluating the performance of Algorithm 1 in relation to the theoretical results explored at the beginning of Section 5. This analysis further supports the choice of the tolerance selected in the following tests. Specifically, we fix $2000$ iterations to be executed, set the time step $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$ and implement the algorithm for several refinements of the space grid. The results are illustrated in Figure 2 and summarized in Table 1. In Figure 2 (left), the behavior of $(\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k})/2$ is displayed for several values of the spatial step $\Delta x$. As expected, this quantity decreases as $k$ increases. Consequently, setting a small tolerance in the stopping condition requires the algorithm to perform a large number of iterations. In Figure 2 (right), the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is shown for several values of $\Delta x$. Consistently with the discussion at the beginning of Section 5, this error stabilizes and reaches a constant value after a sufficient number of iterations. This result suggests that, for a fixed $\Delta x$, running all $2000$ iterations is unnecessary; instead, the algorithm can be stopped once the plateau is reached. Examining the behavior as $\Delta x$ varies, as once again predicted by the previous theoretical analysis, it can be seen that the value of the aforementioned plateau decreases as $\Delta x$ decreases. Additionally, the number of iterations required to reach the plateau increases (or equivalently, the tolerance $\varepsilon$ required to stop the algorithm must be smaller) as $\Delta x$ decreases. These observations justify the following choice of tolerance proportional to $\Delta x$.
We now implement Algorithm 1 with $\varepsilon=\Delta x/10$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$ and several refinements of the space grid. The results are presented in columns 8–10 of Table 2. Column 10 shows a decreasing trend in the error as $\Delta x$ decreases. As previously justified theoretically and observed in Figure 2 (right), this trend is independent of the choice of tolerance scaled by $\Delta x$. Indeed, comparing the errors in column 4 of Table 1 with those in column 10 of Table 2, the approximated critical value is computed with the same order of error. Additionally, we implement Algorithm 1 using the same tolerance $\varepsilon=\Delta x/10$ but varying the time step: $\Delta t=\Delta x^{5/6}$ (columns 2–4 of Table 2) and $\Delta t=\Delta x/2$ (columns 5–7 of Table 2). Table 2 reveals that the worst approximation is obtained with $\Delta t=\Delta x^{5/6}$, while choosing $\Delta t=\Delta x/2$ produces errors comparable to those achieved with $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma}/12$ preserving almost the same number of iterations. This choice offers the benefit of a larger time step, thereby reducing the computational cost.
Finally, we evaluate Algorithm 2 with $\varepsilon=\Delta x/10$ and report the results in Table 3. Columns 2–4 correspond to $\Delta t=\Delta x^{5/6}$, columns 5–7 to $\Delta t=\Delta x/2$, and columns 8–10 to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$. Comparing this table with Table 2, we find that Algorithm 2 exhibits a similar approximation to Algorithm 1 and reduces the number of iteration on average by $67\%$ for $\Delta t=\Delta x^{5/6}$, $80\%$ for $\Delta t=\Delta x/2$ and $83\%$ for $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$.
Figure 2: Algorithm 1. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$, $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x/12$ and $2000$ fixed iterations.
@@
Table 1: Algorithm 1. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/12$ and $2000$ fixed iterations.
@@
$2.00\cdot 10^{-1}$ $40$ $1.91$ $9.11\cdot 10^{-1}$ $19$ $1.20$ $1.95\cdot 10^{-1}$ $18$ $1.16$ $1.63\cdot 10^{-1}$
@@
Table 2: Algorithm 1. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/12$.
@@
$2.00\cdot 10^{-1}$ $4$ $1.91$ $9.13\cdot 10^{-1}$ $7$ $1.18$ $1.79\cdot 10^{-1}$ $5$ $1.15$ $1.50\cdot 10^{-1}$
$1.00\cdot 10^{-1}$ $12$ $1.29$ $2.92\cdot 10^{-1}$ $10$ $1.07$ $7.09\cdot 10^{-2}$ $8$ $1.06$ $6.32\cdot 10^{-2}$
@@
$1.25\cdot 10^{-2}$ $65$ $1.02$ $1.59\cdot 10^{-2}$ $6$ $1.00$ $5.98\cdot 10^{-4}$ $7$ $1.00$ $5.06\cdot 10^{-4}$
Table 3: Algorithm 2. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/12$.
@@ 5.1.2 Hamiltonian Independent of $s$
We begin by performing a preliminary analysis of Algorithm 1 through two tests, both using $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$. The first test employs $\varepsilon=\Delta x/10$ (columns 2–4 of Table 4), while the second test fixes $2000$ iterations to be executed (Figure 3 and columns 5–7 of Table 4). As observed in the previous problem, the decreasing behavior of $(\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k})/2$, shown in Figure 3 (left), confirms that a smaller tolerance $\varepsilon$ in Algorithm 1 leads the algorithm to perform a larger number of iterations before stopping. However, unlike the previous problem where the Hamiltonians depend on $s$, the current tests do not approximate critical values with error of the same order of magnitude. Indeed, as illustrated in Figure 3 (right), the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ decreases as $k$ increases, without reaching a plateau. In this case, the decreasing trend of errors as $\Delta x$ decreases, presented in column 4 of Table 4, dependents on the tolerance proportional to $\Delta x$. This is because as $\Delta x$ decreases, the tolerance $\varepsilon$ also decreases, requiring to an increase in the number of iterations (see column 2 of Table 4). What has just been observed does not contradict the theoretical analysis conducted at the beginning of Section 5. Indeed, as derived from Theorem 4.4 (or equivalently, from Proposition 4.6), the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is bounded above by a term inversely proportional to $k$ (or equivalently, by $(\overline{c}_{k}^{\Delta}-\underline{c}_{k}^{\Delta})/2$) and an additional term depending on $\min\left\{\frac{\Delta x}{\sqrt{\Delta t}},\frac{\Delta x^{2}}{\Delta t^{3/2}}\right\}$, independent of $k$. The latter term originates exclusively from the approximation error estimate for the semi-Lagrangian scheme applied to ($\mathcal{H}$JE). However, as thoroughly analyzed in [CarliniCoscettiPozza24], when the Hamiltonians are independent of $s$ and condition (14) holds, the approximation error is primarily due to the approximation of the minimum in the scheme. As a result, this error is much smaller than the estimate in Theorem 4.1. Consequently, in the error estimate for $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$, the dominant term is that one proportional to $1/k$ (or equivalently, the dominant term is $(\overline{c}_{k}^{\Delta}-\underline{c}_{k}^{\Delta})/2$), which explains the absence of a plateau in Figure 3 (right).
Next, we implement Algorithm 1 with the same time step $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$, but with a smaller tolerance $\varepsilon=\Delta x/100$ (columns 8–11 in Table 5). This yields smaller errors compared to the previous test shown in column 4 of Table 4. As just discussed, this reduction in error can be attributed to the choice of a smaller tolerance. Additionally, the error exhibits a decreasing trend as $\Delta x$ decreases, which is still caused by the tolerance being proportional to the spatial step.
Furthermore, we apply Algorithm 1 keeping $\varepsilon=\Delta x/100$ but exploring different time step choices: $\Delta t=\Delta x^{5/6}$ (columns 2–4 of Table 5) and $\Delta t=\Delta x/2$ (columns 5–7 of Table 5). It can be seen that the largest error occurs when $\Delta t=\Delta x^{5/6}$. On the other hand, setting $\Delta t=\Delta x/2$ results in errors of a comparable magnitude while preserving an almost identical number of iterations. This latter choice is particularly advantageous, as the larger time step leads to a reduction in computational cost.
Finally, we implement Algorithm 2 with $\varepsilon=\Delta x/100$ and report the results in Table 6, whose columns 2–4 refer to $\Delta t=(\Delta x)^{5/6}$, columns 5–7 to $\Delta t=\Delta x/2$, and columns 8–10 to $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma}/9.1$. A comparison with Table 5 reveals that Algorithm 2 achieves a better approximation with substantially fewer iterations. More specifically, the number of iterations is reduced on average by $79\%$ for $\Delta t=(\Delta x)^{5/6}$, $93\%$ for $\Delta t=\Delta x/2$ and $97\%$ for $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma}/9.1$.
@@
Figure 3: Algorithm 1. Test on the triangle with Hamiltonian independent of $s$, $A=1$, $B=0$, $C=1$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$ and $2000$ fixed iterations.
@@
Table 4: Algorithm 1. Test on the triangle with Hamiltonian independent of $s$, $A=1$, $B=0$, $C=1$ and $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$. Columns 2–4 refer to tests with $\varepsilon=\Delta x/10$ and columns 5–7 refer to tests with $2000$ fixed iterations.
@@
Table 5: Algorithm 1. Test on the triangle with Hamiltonian independent on $s$, $A=1$, $B=0$, $C=1$ and $\varepsilon=\Delta x/100$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$.
@@
$2.00\cdot 10^{-1}$ $49$ $0.95$ $5.11\cdot 10^{-2}$ $18$ $1.00$ $1.41\cdot 10^{-3}$ $9$ $1.00$ $5.11\cdot 10^{-5}$
@@
Table 6: Algorithm 2. Test on the triangle with Hamiltonian independent on $s$, $A=1$, $B=0$, $C=1$ and $\varepsilon=\Delta x/100$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$.

### v2

@@ 5.1 Test on a Triangle
plus the reversed arcs (Figure˜1).
[…]
@@ 5.1.1 Hamiltonian Dependent on $s$
We begin evaluating the performance of Algorithm˜1 in relation to the theoretical results explored at the beginning of Section˜5. This analysis further supports the choice of the tolerance selected in the following tests. Specifically, we fix $2000$ iterations to be executed, set the time step $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$ and implement the algorithm for several refinements of the space grid. The results are illustrated in Figure˜2 and summarized in Table˜1. In Figure˜2 (left), the behavior of $(\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k})/2$ is displayed for several values of the spatial step $\Delta x$. As expected, this quantity decreases as $k$ increases. Consequently, setting a small tolerance in the stopping condition requires the algorithm to perform a large number of iterations. In Figure˜2 (right), the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is shown for several values of $\Delta x$. Consistently with the discussion at the beginning of Section˜5, this error stabilizes and reaches a constant value after a sufficient number of iterations. This result suggests that, for a fixed $\Delta x$, running all $2000$ iterations is unnecessary and the algorithm can be stopped once the plateau is reached. Examining the behavior as $\Delta x$ varies, as once again predicted by the previous theoretical analysis, it can be seen that the value of the aforementioned plateau decreases as $\Delta x$ decreases. Additionally, the number of iterations required to reach the plateau increases (or equivalently, the tolerance $\varepsilon$ required to stop the algorithm must be smaller) as $\Delta x$ decreases. These observations justify the following choice of tolerance proportional to $\Delta x$.
Figure 2: Algorithm˜1. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$, $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x/12$ and $2000$ fixed iterations.
@@
Table 1: Algorithm˜1. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/12$ and $2000$ fixed iterations.
We now implement Algorithm˜1 with $\varepsilon=\Delta x/10$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$ and several refinements of the space grid. The results are presented in columns 8–10 of Table˜2. Column 10 shows a decreasing trend in the error as $\Delta x$ decreases. As previously justified theoretically and observed in Figure˜2 (right), this trend is independent of the choice of tolerance scaled by $\Delta x$. Indeed, comparing the errors in column 4 of Table˜1 with those in column 10 of Table˜2, the approximated critical value is computed with the same order of error. Additionally, we implement Algorithm˜1 using the same tolerance $\varepsilon=\Delta x/10$ but varying the time step: $\Delta t=\Delta x^{5/6}$ (columns 2–4 of Table˜2) and $\Delta t=\Delta x/2$ (columns 5–7 of Table˜2). Table˜2 reveals that the worst approximation is obtained with $\Delta t=\Delta x^{5/6}$, while choosing $\Delta t=\Delta x/2$ produces errors comparable to those achieved with $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma}/12$ preserving almost the same number of iterations. This choice offers the benefit of a larger time step, thereby reducing the computational cost.
@@
$2.00\cdot 10^{-1}$ $40$ $1.91$ $1\cdot 10^{-1} 1919$ $1.20$ $1.95\cdot 10^{-1}$ $18$ $1.16$ $1.63\cdot 10^{-1}$
@@
Table 2: Algorithm˜1. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/12$.
Finally, we evaluate Algorithm˜2 with $\varepsilon=\Delta x/10$ and report the results in Table˜3. Columns 2–4 correspond to $\Delta t=\Delta x^{5/6}$, columns 5–7 to $\Delta t=\Delta x/2$, and columns 8–10 to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$. Comparing this table with Table˜2, we find that Algorithm˜2 exhibits a similar approximation to Algorithm˜1 and reduces the number of iterations on average by $67\%$ for $\Delta t=\Delta x^{5/6}$, $80\%$ for $\Delta t=\Delta x/2$ and $83\%$ for $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/12$.
@@
$2.00\cdot 10^{-1}$ 44 $1.91$ $9.13\cdot 10^{-1}$ 77 $1.18$ $1.79\cdot 10^{-1}$ 55 $1.15$ $1.50\cdot 10^{-1}$
$1.00\cdot 10^{-1}$ $12$ $1.29$ $2.92\cdot 10^{-1}$ $10$ $1.07$ $7.09\cdot 10^{-2}$ 88 $1.06$ $6.32\cdot 10^{-2}$
@@
$1.25\cdot 10^{-2}$ $65$ $1.02$ $1.59\cdot 10^{-2}$ 66 $1.00$ $5.98\cdot 10^{-4}$ 77 $1.00$ $5.06\cdot 10^{-4}$
Table 3: Algorithm˜2. Test on the triangle with Hamiltonian dependent on $s$, $A=2/3$, $B=1/2$, $C=1$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/12$.
@@ 5.1.2 Hamiltonian Independent of $s$
We begin by performing a preliminary analysis of Algorithm˜1 through two tests, both using $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$. The first test employs $\varepsilon=\Delta x/10$ (columns 2–4 of Table˜4), while the second test fixes $2000$ iterations to be executed (Figure˜3 and columns 5–7 of Table˜4). As observed in the previous problem, the decreasing behavior of $(\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k})/2$, shown in Figure˜3 (left), confirms that a smaller tolerance $\varepsilon$ in Algorithm˜1 leads the algorithm to perform a larger number of iterations before stopping. However, unlike the previous problem where the Hamiltonians depend on $s$, the current tests do not approximate critical values with error of the same order of magnitude. Indeed, as illustrated in Figure˜3 (right), the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ decreases as $k$ increases, without reaching a plateau. In this case, the decreasing trend of errors as $\Delta x$ decreases, presented in column 4 of Table˜4, dependents on the tolerance proportional to $\Delta x$. This is because as $\Delta x$ decreases, the tolerance $\varepsilon$ also decreases, requiring to an increase in the number of iterations (see column 2 of Table˜4). What has just been observed does not contradict the theoretical analysis conducted at the beginning of Section˜5. Indeed, as derived from Theorem˜4.4 (or equivalently, from Proposition˜4.5), the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is bounded above by a term inversely proportional to $k$ (or equivalently, by $(\overline{c}_{k}^{\Delta}-\underline{c}_{k}^{\Delta})/2$) and an additional term depending on $\min\left\{\frac{\Delta x}{\sqrt{\Delta t}},\frac{\Delta x^{2}}{\Delta t^{3/2}}\right\}$, independent of $k$. The latter term originates exclusively from the approximation error estimate for the semi-Lagrangian scheme applied to ($\mathcal{H}$JE). However, as thoroughly analyzed in [11], when the Hamiltonians are independent of $s$ and condition (14) holds, the approximation error is primarily due to the approximation of the minimum in the scheme. As a result, this error is much smaller than the estimate in Theorem˜4.1. Consequently, in the error estimate for $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$, the dominant term is that one proportional to $1/k$ (or equivalently, the dominant term is $(\overline{c}_{k}^{\Delta}-\underline{c}_{k}^{\Delta})/2$), which explains the absence of a plateau in Figure˜3 (right).
@@
Figure 3: Algorithm˜1. Test on the triangle with Hamiltonian independent of $s$, $A=1$, $B=0$, $C=1$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$ and $2000$ fixed iterations.
@@
Table 4: Algorithm˜1. Test on the triangle with Hamiltonian independent of $s$, $A=1$, $B=0$, $C=1$ and $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$. Columns 2–4 refer to tests with $\varepsilon=\Delta x/10$ and columns 5–7 refer to tests with $2000$ fixed iterations.
Next, we implement Algorithm˜1 with the same time step $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$, but with a smaller tolerance $\varepsilon=\Delta x/100$ (columns 8–11 in Table˜5). This yields smaller errors compared to the previous test shown in column 4 of Table˜4. As just discussed, this reduction in error can be attributed to the choice of a smaller tolerance. Additionally, the error exhibits a decreasing trend as $\Delta x$ decreases, which is still caused by the tolerance being proportional to the spatial step.
Furthermore, we apply Algorithm˜1 keeping $\varepsilon=\Delta x/100$ but exploring different time step choices: $\Delta t=\Delta x^{5/6}$ (columns 2–4 of Table˜5) and $\Delta t=\Delta x/2$ (columns 5–7 of Table˜5). It can be seen that the largest error occurs when $\Delta t=\Delta x^{5/6}$. On the other hand, setting $\Delta t=\Delta x/2$ results in errors of a comparable magnitude while preserving an almost identical number of iterations. This latter choice is particularly advantageous, as the larger time step leads to a reduction in computational cost.
@@
Table 5: Algorithm˜1. Test on the triangle with Hamiltonian independent on $s$, $A=1$, $B=0$, $C=1$ and $\varepsilon=\Delta x/100$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$.
Finally, we implement Algorithm˜2 with $\varepsilon=\Delta x/100$ and report the results in Table˜6, whose columns 2–4 refer to $\Delta t=(\Delta x)^{5/6}$, columns 5–7 to $\Delta t=\Delta x/2$, and columns 8–10 to $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma}/9.1$. A comparison with Table˜5 reveals that Algorithm˜2 achieves a better approximation with substantially fewer iterations. More specifically, the number of iterations is reduced on average by $79\%$ for $\Delta t=(\Delta x)^{5/6}$, $93\%$ for $\Delta t=\Delta x/2$ and $97\%$ for $\Delta t=\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma}/9.1$.
@@
$2.00\cdot 10^{-1}$ $49$ $0.95$ $5.11\cdot 10^{-2}$ $18$ $1.00$ $1.41\cdot 10^{-3}$ 99 $1.00$ $5.11\cdot 10^{-5}$
@@
Table 6: Algorithm˜2. Test on the triangle with Hamiltonian independent on $s$, $A=1$, $B=0$, $C=1$ and $\varepsilon=\Delta x/100$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.1$.